### v3

@@ Abstract
Embodied artificial intelligence (EAI) integrates advanced AI models into physical entities for real-world interaction. The emergence of foundation models as the “brain” of EAI agents for high-level task planning has shown promising results. However, the deployment of these agents in physical environments presents significant safety challenges. For instance, a housekeeping robot lacking sufficient risk awareness might place a metal container in a microwave, potentially causing a fire. To address these critical safety concerns, comprehensive pre-deployment risk assessments are imperative. This study introduces EAIRiskBench , a novel framework for automated physical risk assessment in EAI scenarios. EAIRiskBench   employs a multi-agent cooperative system that leverages various foundation models to generate safety guidelines, create risk-prone scenarios, make task planning, and evaluate safety systematically. Utilizing this framework, we construct EAIRiskDataset , comprising diverse test cases across various domains, encompassing both textual and visual scenarios. Our comprehensive evaluation of state-of-the-art foundation models reveals alarming results: all models exhibit high task risk rates (TRR), with an average of 95.75% across all evaluated models. Notably, even GPT-4o, widely regarded as one of the most advanced models, demonstrates a TRR of 94.03%, underscoring the pervasive lack of risk identification and avoidance capabilities in complex physical environments. To address these challenges, we further propose two prompting-based risk mitigation strategies. While these strategies demonstrate some efficacy in reducing TRR, the improvements are limited, still indicating substantial safety concerns. This study provides the first large-scale assessment of physical risk awareness in EAI agents. Our findings underscore the critical need for enhanced safety measures in EAI systems and provide valuable insights for future research directions in developing safer embodied artificial intelligence system.
@@
Embodied Artificial Intelligence, Safety Assessment, Foundation Models
@@ 2.3 Evaluation of Foundation Models as “Brain” of EAI Agents
Based on the proposed EAIRiskBench and EAIRiskDataset, we conduct extensive evaluation encompassing a diverse range of models. Our study included both unimodal text-only large language models (LLMs) and multimodal vision-language models (VLMs), covering open-source and closed-source variants across various scales. Both unimodal text-only LLMs and multimodal VLMs are evaluated on the textual scenarios of the EAIRiskDataset, while the multimodal VLMs are further assessed on the visual scenarios. To quantify the performance, we introduced two key metrics:
  • Task Risk Rate (TRR): This metric represents the percentage of instances where the generated high-level plans contain potential risks or unsafe actions. A higher TRR indicates poorer risk perception and avoidance capabilities.
  • Task Effectiveness Rate (TER): This measures the percentage of instances in which the generated high-level plans are executable based on the given skill set. A higher TER suggests better task completion capability.
@@ 4.2 Risky Scene Generation
  • Object positions: The object positions determine spatial arrangement of objects in the scene. We use a triple <<<object_1, relation, object_2>>> to represent relative positions, where relation is spatial descriptor such as above, below, behind. These relational positions help simulate realistic scenes by considering how objects are typically organized and interact with each other in an environment, allowing the constructed scenes to more accurately reflect possible real-world configurations.
@@ 4.5 Evaluated Foundation Models
5 Data availability
The EAIRiskDataset used in this study is publicly available at https://github.com/zihao-ai/EAIRiskBench. This repository contains the complete dataset used for evaluating the risk awareness capabilities of EAI agents, including both textual and visual scenarios. All datasets are provided under the MIT license.
6 Code availability
The complete codebase for EAIRiskBench is open-source and can be accessed at https://github.com/zihao-ai/EAIRiskBench. This repository provides all necessary code to reproduce our results and extend our work for future research in embodied artificial intelligence. All source code is provided under the MIT license.

### v4

@@ Abstract
Embodied artificial intelligence (EAI) integrates advanced AI models into physical entities for real-world interaction. The emergence of foundation models as the “brain” of EAI agents for high-level task planning has shown promising results. However, the deployment of these agents in physical environments presents significant safety challenges. For instance, a housekeeping robot lacking sufficient risk awareness might place a metal container in a microwave, potentially causing a fire. To address these critical safety concerns, comprehensive pre-deployment risk assessments are imperative. This study introduces EAIRiskBench, a novel framework for automated physical risk assessment in EAI scenarios. EAIRiskBench employs a multi-agent cooperative system that leverages various foundation models to generate safety guidelines, create risk-prone scenarios, make task planning, and evaluate safety systematically. Utilizing this framework, we construct EAIRiskDataset, comprising diverse test cases across various domains, encompassing both textual and visual scenarios. Our comprehensive evaluation of state-of-the-art foundation models reveals alarming results: all models exhibit high task risk rates (TRR), with an average of 95.75% across all evaluated models. Notably, even GPT-4o, widely regarded as one of the most advanced models, demonstrates a TRR of 94.03%, underscoring the pervasive lack of risk identification and avoidance capabilities in complex physical environments. To address these challenges, we further propose two prompting-based risk mitigation strategies. While these strategies demonstrate some efficacy in reducing TRR, the improvements are limited, still indicating substantial safety concerns. This study provides the first large-scale assessment of physical risk awareness in EAI agents. Our findings underscore the critical need for enhanced safety measures in EAI systems and provide valuable insights for future research directions in developing safer embodied artificial intelligence system. Data and code are available at https://github.com/zihao-ai/EAIRiskBench.
@@
Embodied artificial intelligence, AI safety, Foundation models, Large language models, AI agents
@@ 2.3 Evaluation of Foundation Models as “Brain” of EAI Agents
Based on the proposed EAIRiskBench and EAIRiskDataset, we conduct extensive evaluation encompassing a diverse range of models. Our study included both unimodal text-only large language models (LLMs) and multimodal vision-language models (VLMs), covering open-source and closed-source variants across various scales. Both unimodal text-only LLMs and multimodal VLMs are evaluated on the textual scenarios of the EAIRiskDataset, while the multimodal VLMs are further assessed on the visual scenarios.
To quantify the performance of EAI agents in terms of safety and effectiveness, we introduce two key metrics: Task Risk Rate (TRR) and Task Effectiveness Rate (TER). Let $\mathcal{T}=\{t_{1},...,t_{N}\}$ denote a set of $N$ safety tips, each $t_{i}$ associated with corresponding observation, and task instruction. For each case, an EAI agent generates a high-level plan $\bm{p}_{i}=[s_{1},...,s_{k_{i}}]$ consisting of $k_{i}$ steps. We define two binary indicator functions: safety indicator $\mathbb{I}_{s}(\bm{p}_{i},s_{i})=1$ when $\bm{p}_{i}$ contains potential risk denoted in $s_{i}$; effectiveness indicator $\mathbb{I}_{e}(p_{i})=1$ for any step $s_{i}\in\bm{p}_{i}$ is executable. TRR, representing the percentage of plans containing potential risks, is defined as:
$\text{TRR}=\frac{\sum_{i=1}^{N}\mathbb{I}_{s}(\bm{p}_{i},s_{i})}{N}.$ (1)
TER, measuring the percentage of executable plans, is defined as:
$\text{TER}=\frac{\sum_{i=1}^{N}\mathbb{I}_{e}(\bm{p}_{i})}{N}.$ (2)
Higher TRR indicates poorer risk identification and avoidance capabilities, while higher TER suggests better task completion capability.
@@ 4.2 Risky Scene Generation
  • Object positions: The object positions determine spatial arrangement of objects in the scene. We use a triple $<$object_1, relation, object_2$>$ to represent relative positions, where relation is spatial descriptor such as above, below, behind. These relational positions help simulate realistic scenes by considering how objects are typically organized and interact with each other in an environment, allowing the constructed scenes to more accurately reflect possible real-world configurations.